# DaWin: Training-free Dynamic Weight Interpolation for Robust Adaptation

### v1

@@ 2.2 Pilot Study
Our goal. We begin by conducting a pilot study to understand the benefits of dynamic interpolation, which adapts interpolation coefficients at a finer granularity (such as sample level, $\lambda(x)$), as opposed to using global coefficient $\lambda$ for all samples (Wortsman et al., 2022b; a). To explore the upper limits of these methods, we experiment with ground truth labels as an oracle to estimate upper-bound performance and understand the maximum potential of model interpolation methods. This study will further guide our method design in Sec. 3 by approximating the oracle performance.
@@ 3.2 Dynamic Interpolation via Entropy Ratio
Here, $H$ represents the entropy over the output probability distribution of model $f(\cdot;\theta)$ given input $x$. The value of $\lambda(x)$ approaches 1 when model $f(\cdot;\theta_{1})$ exhibits lower entropy (greater expertise), whereas it approaches 0 if the entropy of model $f(\cdot;\theta_{1})$ becomes higher. Figure 3 illustrates the overall framework. Intuitively, this approach is analogous to classifier selection methods (Giacinto & Roli, 2001; Ko et al., 2008), which aims to dynamically select suitable classifier(s) given a test sample. However, DaWin differs in terms of its motivation, expertise metric, and goal as we pursue finding the interpolation coefficients that may induce better performance compared with the best model selection method (See Table 5). We further discuss the connection between DaWin and the selection-based method in Sec. 6. Note that our sample-wise expertise estimations do not require any hyperparameters for computing interpolation coefficients. This eliminates the need for intensive hyperparameter tuning often required in prior works (Wortsman et al., 2022b; Ilharco et al., 2023) to achieve the best result. Meanwhile, if we assume access to all test samples simultaneously, we can refine the coefficient computation by introducing per-domain expertise offset terms (please see Appendix A.1). We use this offset adjustment by default, which brings consistent improvements.
Sample-wise entropy valley hypothesis. One of the most popular hypotheses that explain the success of weight interpolations is linear mode connectivity (LMC; Frankle et al. (2020)), which ensures that interpolated models are also laid on good solution space as well as low-loss converged individual models. Although entropy is a concave function that denies the LMC in nature, we believe that a similar statement can be claimed, which we refer to sample-wise entropy valley as follows:
@@
As empirical validation of this hypothesis, Figure 6 shows that the interpolated model actually induces smaller entropy than individuals (i) on average, and DaWin achieves lower entropy than static interpolation (ii). Given the strong correlation between entropy and X-entropy, this hypothesis advocates the benefit of DaWin than static interpolation methods.
@@ 4.3 Results on Multi-task Learning
Table 6: Multi-task learning performance. We use CLIP ViT-B/32 for evaluation across eight benchmark tasks. $\dagger$ denotes optimal-like performance using the ground truth domain indicator to select models for each domain. We report the layer-wise method for AdaMerging (Yang et al., 2024b) here, which surpasses those of the task-wise method. Note that AdaMerging and Pareto Merging both require a non-trivial amount of training.
@@ 6 Discussion
Suppose we have $M$ different models $\{f(\cdot;\theta_{j})\}_{j=1}^{M}$ parameterized by $\{\theta_{j}\}_{j=1}^{M}$ with a homogeneous architecture defined by $f(\cdot)$. Let $\lambda(x)=(\lambda_{1}(x),\dots,\lambda_{M}(x))$ be the sample-wise interpolation coefficient vector given $x$. If the models that render a correct prediction for $x$ always have smaller output entropy $H(\cdot)$ compared with the models rendering an incorrect prediction, DaWin’s interpolation coefficients of true experts for the sample $x$ are always greater than $\frac{1}{M}$. That is,
$\lambda_{j\in\mathcal{J}}(x)\geq\frac{1}{M}\;\;$ $\text{where}\;\;\mathcal{J}=\{i|\operatorname*{arg\,max}_{c}[f(x;\theta_{i})]_{c}=y\}\;\;\text{if}\;\;H(f(x;\theta_{j\in\mathcal{J}}))\leq H(f(x;\theta_{k\notin\mathcal{J}}))\;\;\forall j\;\text{and}\;k,$
where $[f(x;\theta)]_{c}$ indicates the probability mass for class $c$. Lemma 6.1 implies that, under the entropy-dominancy assumption, DaWin always produces per-sample expert-biased coefficient vectors, which result in the interpolated models being biased towards true experts, i.e., models that produce correct prediction given $x$. This desirable behavior is aligned with the motivation of dynamic classifier selection discussed in Sec. 4.2, whereas DaWin conducts interpolation rather than selection. Meanwhile, as the number of models participating in interpolation increased, samples that at least one model correctly classifies also increased. Therefore, the coverage of DaWin for weighing the correct experts expands accordingly. This analysis endows a potential clue for remarkable gains observed in the multi-task setting, which conducts merging beyond two models (See Sec. 4.3).
@@ Acknowledgments
We thank the researchers and interns at NAVER AI Lab – Byeongho Heo, Taekyung Kim, Sanghyuk Chun, Sehyun Kwon, Yong-Hyun Park, and Jaeyoo Park – for their invaluable feedback. Most of the experiments are based on the NAVER Smart Machine Learning (NSML) platform (Sung et al., 2017). We also appreciate on the constructive comments from Max Khanov, Hyeong Kyu Choi, and Seongheon Park at University of Wisconsin–Madison. This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2022-00143911, AI Excellence Global Innovative Leader Education Program) and the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT)(RS-2024-00457216, 2022R1A4A3033874)
@@ A.1 Algorithm and additional details on DaWin
While the eq equation 1 yields sample-wise interpolation coefficients solely based on the per-sample expertise estimation, if we can access the whole test samples simultaneously, we can refine the coefficient computation by introducing domain-wise offset terms. These terms are automatically estimated by the per-domain entropy of each model. Given the significant performance improvements with domain-wise dynamic interpolation (c.f. Table 1), we expect domain-wise expertise per model can be a complimentary benefit to compute the interpolation coefficient. Therefore, we modify the sample-wise coefficient term in Eq. 1 as below:
@@ A.2.3 Implementation Details
On DaWin’s evaluation, we first get the entropy of batch test samples from the interpolation candidate models [^5] Candidate models are constituted with the pre-trained and fine-tuned models for robust fine-tuning setting, the task-specific eight fine-tuned models for multi-task learning setting. (wherein temperature scaling (Guo et al., 2017) is applied in the robust fine-tuning setup with ID validation set), then we compute interpolation coefficients by building a softmax-like model expertise ratio term with exponentiated negative entropy of each model. We further perform the mixture modeling over the batch test samples and finally conduct dynamic model merging with interpolation coefficients corresponding to estimated membership from the fitted mixture model to obtain the prediction per sample. About the Beta (on robust fine-tuning setup) and Dirichlet (on multi-tasks learning setup) mixture modeling on interpolation coefficients for DaWin, we set $K$ to 3, 5, 2 for ViT-{B/32, B/16, L/14} in the robust fine-tuning and $K=1$ in the multi-task setups. Unless otherwise mentioned, we adopt the offset adjustment term (Eq. equation 2) by default, assuming that the entire test samples per task are available and fit the Beta (and Dirichlet) mixture model on the entire coefficients per task, likewise assumption of Yang et al. (2024b; a); Chen & Kwok (2024).
@@ A.4 Missing proof
Suppose we have $M$ different models $\{f(\cdot;\theta_{j})\}_{j=1}^{M}$ parameterized by $\{\theta_{j}\}_{j=1}^{M}$ with a homogeneous architecture defined by $f(\cdot)$. Let $\lambda(x)=(\lambda_{1}(x),\dots,\lambda_{M}(x))$ be the sample-wise interpolation coefficient vector given $x$. If the models that render a correct prediction for $x$ always have smaller output entropy $H(\cdot)$ compared with the models rendering an incorrect prediction, DaWin’s interpolation coefficients of true experts for the sample $x$ has always greater than $\frac{1}{M}$. That is
$\lambda_{j\in\mathcal{J}}(x)\geq\frac{1}{M}\;\quad$ $\text{where}\;\;\mathcal{J}=\{i|\operatorname*{arg\,max}_{c}[f(x;\theta_{i})]_{c}=y\}$
$\text{if}\;\;H(f(x;\theta_{j\in\mathcal{J}}))\leq H(f(x;\theta_{k\notin\mathcal{J}}))\;\;\text{for all}\;\;j\;\text{and}\;k.$
where $[f(x;\theta)]_{c}$ indicates the probability mass for class $c$.

### v2

@@ 2.2 Pilot Study
Our goal. We begin by conducting a pilot study to understand the benefits of dynamic interpolation, which adapts interpolation coefficients at a finer granularity (such as sample level, $\lambda(x)$), as opposed to using global coefficient $\lambda$ for all samples (Wortsman et al., 2022b; a). To explore the upper limits of these methods, we experiment with ground truth labels as an oracle [^1] We use the term oracle to denote a scenario where we utilize the ground truth test label, and it should not be confused with optimal. Optimality of $\lambda^{*}(x)$ regarding weight interpolation is further discussed in Tab C. to estimate upper-bound performance and understand the maximum potential of model interpolation methods. This study will further guide our method design in Sec. 3 by approximating the oracle performance.
@@ 3.2 Dynamic Interpolation via Entropy Ratio
Here, $H$ represents the entropy over the output probability distribution of model $f(\cdot;\theta)$ given input $x$. The value of $\lambda(x)$ approaches 1 when model $f(\cdot;\theta_{1})$ exhibits lower entropy (greater expertise), whereas it approaches 0 if the entropy of model $f(\cdot;\theta_{1})$ becomes higher. Figure 3 illustrates the overall framework. Intuitively, this approach is analogous to classifier selection methods (Giacinto & Roli, 2001; Ko et al., 2008), which aims to dynamically select suitable classifier(s) given a test sample. However, DaWin differs in terms of its motivation, expertise metric, and goal as we pursue finding the interpolation coefficients that may induce better performance compared with the best model selection method (See Table 5). We further discuss the connection between DaWin and the selection-based method in Sec. 6. Note that our sample-wise expertise estimations do not require any hyperparameters for computing interpolation coefficients. This eliminates the need for intensive hyperparameter tuning often required in prior works (Wortsman et al., 2022b; Ilharco et al., 2023) to achieve the best result. Meanwhile, if we assume access to all test samples simultaneously, we can refine some potentially unstable coefficient estimations by introducing per-domain expertise offset terms (please see Sec. A.1). We use this offset adjustment by default and utilize the calibrated output of models via temperature scaling to bypass the overconfidence issue (Guo et al., 2017).
Sample-wise entropy valley hypothesis. One of the most popular hypotheses that explain the success of weight interpolations is linear mode connectivity (LMC; Frankle et al. (2020)), which ensures that interpolated models are also laid on good solution space as well as low-loss converged individual models. Although entropy is a concave function for the probability density, we believe that a similar statement can be claimed, which we refer to sample-wise entropy valley as follows:
@@
As empirical validation of this hypothesis, Figure 6 and Figure H show that the interpolated model actually induces smaller entropy than individuals (i) on average, and DaWin achieves lower entropy than static interpolation (ii). Given the strong correlation between entropy and X-entropy, this hypothesis advocates the benefit of DaWin than static interpolation methods.
@@ 4.3 Results on Multi-task Learning
Table 6: Multi-task learning performance. We use CLIP ViT-B/32 for evaluation across eight benchmark tasks. $\dagger$ denotes using the ground truth domain indicator to select expert models for each domain. We report the layer-wise method for AdaMerging (Yang et al., 2024b) here, which surpasses those of the task-wise method. Note that AdaMerging and Pareto Merging both require a non-trivial amount of training.
@@ 6 Discussion
Suppose we have $M$ different models $\{f(\cdot;\theta_{j})\}_{j=1}^{M}$ parameterized by $\{\theta_{j}\}_{j=1}^{M}$ with a homogeneous architecture defined by $f(\cdot)$. Let $\lambda(x)=(\lambda_{1}(x),\dots,\lambda_{M}(x))$ be the sample-wise interpolation coefficient vector given $x$, and $[f(x;\theta)]_{c}$ denotes the probability mass for class $c$. Then, we have
$\lambda_{j\in\mathcal{J}}(x)\geq\frac{1}{M}\;\;\;\;$ $\text{if}\;\;H(f(x;\theta_{j\in\mathcal{J}}))\leq H(f(x;\theta_{k\notin\mathcal{J}}))\;\;\forall j\;\text{and}\;k,$
$\text{where}\;\;\mathcal{J}=\{i|\operatorname*{arg\,max}_{c}[f(x;\theta_{i})]_{c}=y\}.$
Lemma 6.1 implies that, under the entropy-dominancy assumption, DaWin always produces per-sample expert-biased coefficient vectors, which result in the interpolated models being biased towards true experts, i.e., models that produce correct prediction given $x$. This desirable behavior is aligned with the motivation of dynamic classifier selection discussed in Sec. 4.2, whereas DaWin conducts interpolation rather than selection. Meanwhile, as the number of models participating in interpolation increased, samples that at least one model correctly classifies also increased. Therefore, the coverage of DaWin for weighing the correct experts expands accordingly. This analysis endows a potential clue for remarkable gains observed in the multi-task setting, which conducts merging beyond two models (See Sec. 4.3).
@@ Acknowledgments
We thank the researchers and interns at NAVER AI Lab – Byeongho Heo, Taekyung Kim, Sanghyuk Chun, Sehyun Kwon, Yong-Hyun Park, and Jaeyoo Park – for their invaluable feedback. Most of the experiments are based on the NAVER Smart Machine Learning (NSML) platform (Sung et al., 2017). We also appreciate on the constructive comments from Max Khanov, Hyeong Kyu Choi, and Seongheon Park at University of Wisconsin–Madison. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2024-00457216, NRF-RS-2024-00466956, FY2025).
@@ A.1 Algorithm and additional details on DaWin
While Eq. 1 yields sample-wise interpolation coefficients solely based on the per-sample expertise estimation, if we can access the whole test samples simultaneously, we can refine the coefficient computation by introducing domain-wise offset terms. These terms are automatically estimated by the per-domain entropy of each model. Given the significant performance improvements with domain-wise dynamic interpolation (c.f. Table 1), we expect domain-wise expertise per model can be a complimentary benefit to compute the interpolation coefficient. Therefore, we modify the sample-wise coefficient term in Eq. 1 as below:
@@ A.2.3 Implementation Details
On DaWin’s evaluation, we first get the entropy of batch test samples from the interpolation candidate models [^6] Candidate models are constituted with the pre-trained and fine-tuned models for robust fine-tuning setting, the task-specific eight fine-tuned models for multi-task learning setting. (wherein temperature scaling (Guo et al., 2017) is applied in the robust fine-tuning setup with ID validation set), then we compute interpolation coefficients by building a softmax-like model expertise ratio term with exponentiated negative entropy of each model. We further perform the mixture modeling over the batch test samples and finally conduct dynamic model merging with interpolation coefficients corresponding to estimated membership from the fitted mixture model to obtain the prediction per sample. About the Beta (on robust fine-tuning setup) and Dirichlet (on multi-tasks learning setup) mixture modeling on interpolation coefficients for DaWin, we set $K$ to 3, 5, 2 for ViT-{B/32, B/16, L/14} in the robust fine-tuning and $K=1$ in the multi-task setups. Unless otherwise mentioned, we adopt the offset adjustment term (Eq. 2) by default, assuming that the entire test samples per task are available and fit the Beta (and Dirichlet) mixture model on the entire coefficients per task, likewise assumption of Yang et al. (2024b; a); Chen & Kwok (2024).
@@ A.3 Additional empirical results
Figure G: Distance between datasets. We visualize the mean maximum discrepancy (MMD) and Frechet inception distance across eight datasets used in multi-task learning benchmarks in the left and right panels, respectively. To compute MMD and FID, we adopt the pre-trained OpenAI CLIP ViT-B/32 and ImageNet pre-trained Inception V3 models as feature extractors respectively.
In Figure G, we visualize the distance between eight datasets used in the multi-task learning setups. The computed cross-dataset distances are matched with the average coefficients derived from DaWin in Figure 7 to some extent.
Figure H: Empirical validation on the sample-wise entropy valley hypothesis. We visualize the average entropy for each dataset across various methods including single model evaluations (ZS and FT), static weight interpolation (WiSE-FT 0.05, …, 0.95), some dynamic weight interpolation methods ($U(0,1)$, $N(0.5,0.2^{2})$), $N(0.8,0.1^{2})$, Max logit ratio, Confidence ratio, and Confidence difference ratio), and our DaWin method.
Similar to the linear mode connectivity phenomenon that prevails between pre-trained and fine-tuned checkpoints of foundation models (Wortsman et al., 2022b; Ramé et al., 2023) which indicates the interpolations between two checkpoints do not increase the expected loss (e.g., X-entropy) on downstream, in Section 3.2, we hypothesized the expected entropy of model output may also show similar convexity trend over interpolations between pre-trained and fine-tuned model checkpoints. In Figure H, we see that this claim empirically holds for the six datasets including ID (ImageNet) and OOD (its five variances), i.e., there exists an interpolated model that achieves smaller entropy than individual models (ZS and FT). Besides, we further observed that there exists a per-sample dynamic interpolation method that achieves much smaller entropy compared to the static interpolation method. Therefore, our two statements in the sample-wise entropy valley hypothesis show empirical evidence.
Table C: Optimality check for $\lambda^{*}(x)$. We provide mean squared error (MSE) between the optimally estimated per-sample interpolation coefficients and our X-entropy ratio-based estimation in Section 2.2. We estimate the optimal coefficient per sample by conducting a finer-grained grid search for the coefficients, e.g., $\{0.05,0.10,...,0.90,0.95\}$ per sample to minimize the X-entropy loss of interpolated model given a sample $x$. As baseline coefficients, $U(0,1)$ and WiSE-FT constant (0.8) denote per sample random uniform variable and a constant scalar 0.8 for all samples respectively.
- ImageNet ImageNetV2 ImageNetR ImageNetA ImageNetSketch ObjectNet
Std. of optimal coef. 0.2827 0.3068 0.3237 0.3287 0.3413 0.3438
$U(0,1)$ 0.1797 0.1823 0.1869 0.1976 0.1980 0.2013
WiSE-FT constant (0.8) 0.1084 0.1390 0.2081 0.2427 0.2036 0.2131
X-entropy ratio 0.0382 0.1559 0.0479 0.0364 0.0400 0.0440
In Section 2.2, we adopted the negative X-entropy as an oracle model expertise metric and used the exponentiated negative X-entropy ratio as our desired per-sample interpolation coefficients. Although the exponentiated negative X-entropy ratio $\lambda^{*}(x)$ has a nice interpretation connected to the density estimation (Sec A.1), it is not guaranteed that the $\lambda^{*}(x)$’s optimality in the context of weight interpolation. To investigate how close our $\lambda^{*}(x)$ are to the true optimal per-sample interpolation coefficients that minimize the X-entropy losses of interpolation model given samples, we measure mean squared error (MSE) between ours and optimally estimated per-sample coefficients (we estimate these by conducting grid search for interpolation coefficient over $\{0.05,0.10,...,0.90,0.95\}$ per sample). Table C presents the results indicating that our $\lambda^{*}(x)$ are generally much closer to the optimally estimated coefficient compared with random uniform variable or tuned WiSE-FT constant coefficient. Given that the standard deviation of oracle coefficients are about from 0.28 to 0.34, the closeness between optimal and X-entropy ratio-based coefficients is remarkable.
Table D: Alternative per-sample interpolation coefficients. We provide results from some different design choices for per-sample dynamic interpolations (Beta Mixture Modeling with $K$=3 applied for efficient inference) on ImageNet distribution shift benchmarks. Although other data-independent or data-dependent methods struggle to strike a good balance between ID and OOD performances, DaWin shows outstanding performance among considered candidates demonstrating its effectiveness well-grounded with its unique motivation.
Category Coefficient per sample ID Acc OOD Acc Avg.
No interpolation (FT) - 78.4 47.1
Static interpolation 0.8 (WiSE-FT) 79.1 51.0
Data-independent $U(0,1)$ 75.9 52.9
Data-independent $N(0.5,0.2^{2})$ 77.5 54.1
Data-independent $N(0.8,0.1^{2})$ 79.1 51.1
Data-dependent Max Logit Ratio 76.6 54.3
Data-dependent Confidence Diff. Ratio 77.5 54.2
Data-dependent Confidence Ratio 77.5 54.2
Data-dependent Entropy Ratio (ours) 78.7 54.4
In Table D, we provide some alternative design choices for functions that estimate per-sample interpolation coefficients from two categories, data-independent and data-dependent. We see that DaWin outperforms other baseline methods indicating our careful design choice of DaWin motivated by a pilot study has non-trivial benefits rather than other heuristic alternatives.
Table E: Performance of DaWin w/ and w/o confidence calibration on ImageNet (ID) and its variants (OOD).
Method ID Acc OOD Acc
WiSE-FT 79.1 51.0
DaWin (default) 78.7 54.4
DaWin w/ bad calibration 76.6 54.0
Meanwhile, DaWin adopts the entropy of output probability distributions of individual models. As mentioned in Section A.2, we apply temperature scaling to individual merging candidate models to get calibrated outputs from them. The natural question is “how does the uncertainty calibration of individual models affect the final performance of DaWin?”. As shown in Table E, DaWin without temperature scaling (denoted as bad calibration) compromises both ID and OOD performance, so we confirm that reliable expertise estimation is crucial to achieving the full potential of DaWin.
Table F: Runtime, FLOPs, and Peak memory allocation comparison between static (WiSE-FT) and dynamic (DaWin) weight interpolation for evaluation on ImageNet variants. Here, we use ViT-B/32 backbone model on NVIDIA A100 GPU(s). DaWin (parallel) conducts the expertise computation process in parallel (i.e., forward evaluations for pre-trained and fine-tuned models are conducted simultaneously) with multiple GPUs for faster inference.
Method Total Runtime (sec) FLOPs per sample Peak memory (MB) per sample OOD Avg. Acc
WiSE-FT 1040 4.4139 454 51.0
DaWin 1802 13.2452 1338 54.4
DaWin (parallel) 1341 8.8313 1338 54.4
Along with Figure 8, to further investigate the trade-offs between DaWin and a simpler method, we provide evaluation results with additional metrics (FLOPs and peak memory allocation) and variants of DaWin in Table F. Here, DaWin (parallel) denotes the setting where we parallelize the evaluations of individual models during the expertise estimation phase before conducting the interpolation [^7] This treatment is orthogonal to the commonly used distributed data-parallel inference mode and is only applicable to DaWin.. DaWin shows three times FLOPs than WiSE-FT, but the actual runtime is less than twice that of WiSE-FT because DaWin does not require intensive hyperparameter tuning, unlike WiSE-FT. Besides, the computational complexity can be further reduced by adopting the parallel inference pipeline for the model expertise estimation phase of DaWin.
Table G: Fine-grained analysis on interpolation candidates. Model Pool denotes the candidate models (fine-tuned on those datasets) we use to build the interpolated model. We simulate some scenarios where we use relevant expert models only or mixed candidates of experts and non-experts to get the interpolated model. Relevant datasets are determined based on task and semantic similarity between datasets.
Scenario Model Pool Evaluation data Task Arithmetic AdaMerging++ DaWin
Experts only SVHN, MNIST SVHN 87.5 94.1 90.5
Mixed SVHN, MNIST, EuroSAT SVHN 84.4 93.7 93.8
Mixed SVHN, MNIST, EuroSAT, RESISC45 SVHN 81.5 93.4 94.8
Experts only EuroSAT, RESISC45 EuroSAT 96.4 98.1 98.6
Mixed SVHN, MNIST, EuroSAT EuroSAT 86.9 97.7 99.6
Mixed SVHN, MNIST, EuroSAT, RESISC45 EuroSAT 89.6 97.7 99.7
Including our method, the success of model merging methods (also traditional ensemble methods) depends on the relation between the training datasets of merging candidate models and the evaluation dataset. In Table G, we systematically analyze the effect of the relation between train and evaluation datasets on the performance of the merged method by constructing a model pool for merging based on the distance between train datasets (where the candidate models are trained) and test datasets. Please refer to Figure G and Figure 7 to check the quantities of similarity between datasets.
We simulate scenarios where all models are relevant to solve the downstream task, or some models are relevant but others are not to solve the downstream task. Although the baseline methods perform well when we have expert models trained on datasets that are relevant to the evaluation dataset, they suffer from performance degeneration as the non-expert models are included in the merging pool. This indicates that there are some interferences between models (and train corresponding datasets) that hurt post-merging performance if we inappropriately determine the interpolation coefficient. Meanwhile, DaWin takes benefits from even non-expert models by leveraging the entropy ratio-based weighting strategy, and the performance is improved as more models participate in merging.
@@ A.4 Missing proof
Suppose we have $M$ different models $\{f(\cdot;\theta_{j})\}_{j=1}^{M}$ parameterized by $\{\theta_{j}\}_{j=1}^{M}$ with a homogeneous architecture defined by $f(\cdot)$. Let $\lambda(x)=(\lambda_{1}(x),\dots,\lambda_{M}(x))$ be the sample-wise interpolation coefficient vector given $x$, and $[f(x;\theta)]_{c}$ denotes the probability mass for class $c$. Then, we have
$\lambda_{j\in\mathcal{J}}(x)\geq\frac{1}{M}\;\;\;\;$ $\text{if}\;\;H(f(x;\theta_{j\in\mathcal{J}}))\leq H(f(x;\theta_{k\notin\mathcal{J}}))\;\;\forall j\;\text{and}\;k,$
$\text{where}\;\;\mathcal{J}=\{i|\operatorname*{arg\,max}_{c}[f(x;\theta_{i})]_{c}=y\}.$